OpenStudy (anonymous):

Help! At the Student Center cafeteria, 20% of those eating lunch are on a diet and 80% are not. Of those on a diet, 90% select a low-fat dessert, and 20% of those not on a diet select a low-fat dessert. Find the probability that a person who selects a low-fat dessert is on a diet.

OpenStudy (lucaz):

I found 9/17, @sourwing what do you think?

OpenStudy (anonymous):

how did you get that?

OpenStudy (anonymous):

Did you draw a tree diagram?

OpenStudy (lucaz):

diet and low-fat = 20% * 90% = 0.18 no diet and low-fat = 80% * 80% = 0.16 total of low-fat is 0.18+0.16=0.34 P(diet)=0.18/0.34

OpenStudy (anonymous):

Yes I tried drawing the tree diagram but I really don't understand it..

OpenStudy (anonymous):

Thanks @lucaz =)

OpenStudy (anonymous):

so you're good?

OpenStudy (anonymous):

For this question yes,but for the rest of the questions it requires drawing a tree diagram.. Could you explain to me how the tree diagram works? I get confuse on where to place the information that is provided in the question.

OpenStudy (anonymous):



OpenStudy (anonymous):

So you're finding P(D | LFD) yes?

OpenStudy (anonymous):

D = diet LFD = low fat dessert ND = not diet NLFD = not low fat dessert

OpenStudy (anonymous):

using conditional probability formula: P(D | LFD) = P(D and LFD) / P(LFD)

OpenStudy (anonymous):

P(LFD) = P(D and LFD) + P(ND + LFD) = (0.2)(0.9) + (0.8)(0.2) and P(D and LFD) = (0.2)(0.9) so, P(D | LFD) = (0.2)(0.9) / [(0.2)(0.9) + (0.8)(0.2)] = 9/17

OpenStudy (anonymous):

Ok, thank you for explaining =)